The Coffee Shoppe sells a coffee blend made
from two coffees, one costing $5/lb and the other costing
$6/lb. If the blended coffee sells for $5.60/lb, find how
much of each coffee is used to obtain the desired blend.
Assume that the weight of the blended coffee is 100 lb.
Thank you for any help with this.

Answer by ankor@dixie-net.com(22740)

The Coffee Shoppe sells a coffee blend made from two coffees, one costing $5/lb
and the other costing $6/lb.
If the blended coffee sells for $5.60/lb, find how much of each coffee is used to obtain the desired blend.
Assume that the weight of the blended coffee is 100 lb.
:
Let x = amt of $6 coffee required
then
(100-x) = amt o $5 coffee
;
A typical mixture problem
6x + 5(100-x) = 5.6(100)
6x + 500 - 5x = 560
6x - 5x = 560 - 500
x = 60 lb of $6 coffee
I'll let you find how much $5 coffee